# LOLA: LLM-Assisted Online Learning Algorithm for Content Experiments

### v1

@@ 4.1 Bandit Framework and LOLA
Under the current framework, we may have more than two headlines in one test (i.e., $K>2$), rendering the previous LLM-based binary classification models inapplicable. To compare multiple arms, we adapt the previous LLM-based classification models to predict CTR by changing the loss function from log loss to mean squared error (MSE) while keeping the training and fine-tuning procedures unchanged. Using CTR as the outcome measure does not adversely affect decision-making performance. For instance, using linear regression with 256-dimensional text embeddings from OpenAI to predict CTR for significant pairs and labeling the headline with the higher predicted CTR as the winner yields an accuracy of around 75%, which is very close to the 77.56% accuracy obtained with logistic regression. This minor difference is likely due to the inherently skewed nature of the CTR distribution and the sensitivity of MSE to outliers.
@@
Given the above formalization of the decision process and the LLM-based CTR prediction model, the remaining key challenge is to design a decision policy that incorporates these LLM-based CTR predictions. Inspired by the 2-Upper Confidence Bounds (2-UCBs) policy designed by (gur2022adaptive), we propose the following LLM-Assisted 2-Upper Confidence Bounds (LLM-2UCBs) algorithm. Essentially, we treat the LLM’s CTR predictions as auxiliary information prior to the start of online experiments, equating this auxiliary CTR information to additional impressions and click outcomes for each arm.
@@
In the training phase of the LLM-2UCBs algorithm, we can utilize various LLM-based methods, including LoRA fine-tuning using the original text information $x$, and embedding-based CTR prediction models where $x$ represents the embedding vector. The feature vector $x$ can also include additional information, such as user-specific data. However, since Upworthy does not collect user information and ultimately recommends the same headline to all users, incorporating user information into the algorithm is unnecessary in this specific context. Nonetheless, we emphasize that our algorithm is flexible and can be adapted to incorporate user information when it is available.
@@
There are two hyperparameters in the LLM-2UCBs. The upper bound control hyperparameter, $\alpha$, also shows up in the regular UCB algorithm. The hyperparameter $n^{\text{aux}}$ is specific to our proposed algorithm and indicates the equivalent auxiliary samples for initialization. At a high level, we can regard the LLM predictions as prior information or the information flow before the algorithm starts. This prediction is approximately equivalent to initially generating $n^{\text{aux}}$ impressions for each arm and observing their outcomes to obtain the sample average CTR, $\bar{\mu}^{\text{aux}}_{k}$. The hyperparameter fine-tuning process is simple. We can utilize another randomly sampled dataset to test combinations of two hyperparameters and select the combination that yields the highest accumulated rewards.
@@
During the online learning phase, the key step is balancing the exploration and exploitation trade-off. We capture this trade-off using the 2-UCBs rule. Essentially, we have two valid UCBs: $U_{k}^{1}$ is the regular UCB, which uses observed impressions and clicks to construct the mean reward estimator and adds a regular upper bound. The second UCB, $U_{k}^{2}$, incorporates LLM predictions. The calculation of $U_{k}^{2}$ is straightforward. The mean reward component is simply the sample average, taking into account the “auxiliary” samples generated by the LLM before the learning phase begins. These auxiliary samples are imaginary and approximate the richness of information contained in the LLM prediction. The next step is to choose the smaller of the two UCBs, $\min\{U_{k}^{1},U_{k}^{2}\}$, as the final UCB for selecting the arm. Note that if $n^{\text{aux}}=0$, the 2-UCBs rule is the same as the standard UCB rule. Conversely, when $n^{\text{aux}}\rightarrow\infty$, the 2-UCBs rule is equivalent to a pure exploitation policy (a.k.a. greedy policy) using the LLM-based CTR prediction. After playing the arm and observing the outcome, we update the impressions, clicks, and mean reward accordingly, following standard bandit algorithm procedures.
The 2-UCB algorithm offers several advantages. First, its nonparametric nature means it does not assume a linear relationship, making it flexible for various applications. Second, it is easy to implement, requiring the tuning of only two hyperparameters. Third, it can be applied to any LLM-based model. Notice that we do not use the contextual information $x$ during the online learning phase. The primary reason for this is to avoid updating the LLM prediction model during the learning phase, which would require significant engineering effort due to the large-scale nature of training LLMs. However, if engineering costs are not a concern, and for the completeness of this paper, we provide a linear bandit algorithm based on (chu2011contextual), taking into account LLM-based information and user features for personalized recommendations. Despite this, the linear bandit has limitations due to its linear assumption. Therefore, we propose another potential solution based on the FALCON algorithm (simchi2022bypassing). We discuss the pros and cons of these bandit algorithms in Appendix B.
Note that this framework is general and can accommodate any number of headlines and does not rely on binary classification and/or assume that differences between headlines are significant. As such, in the rest of the paper, we utilize the full dataset from Upworthy to test the performance of our proposed algorithm and benchmark approaches (unlike the analysis in Section 3, where we focus only on pairs of headlines with significant differences).
4.2 Benchmarks and Implementation Details in Regret Minimization
In this section, we introduce each of the following algorithms and their implementations in detail, as we will compare them in the next subsection.
@@ 4.2 Benchmarks and Implementation Details in Regret Minimization
    Explore-then-Commit (E&C)
@@
    Pure LLM
@@
    Standard UCB (UCB)
@@
    LLM-Assisted 2-Upper Confidence Bounds (LLM-2UCBs)
At a high level, we randomly split our data into three exclusive folds based on the test ID, ensuring that all headlines in the same test are in the same fold and that the same headline does not appear in different folds. Specifically, we use 70% of the data for training, 10% for fine-tuning, and 20% for testing. All hyperparameters are selected using the fine-tuning data, and the performance of the algorithms is evaluated on the test data. Note that only our proposed LLM-2UCBs and the pure LLM method utilize the training data. In our numerical study, we train the linear regression model with 3,072 dimensions text embeddings from OpenAI as inputs to predict CTRs
Upworthy assigns traffic to headlines with equal probability during the experiment phase, [^24] Upworthy explicitly mentioned they didn’t conduct power analysis before the experiment, and based on the disclosed impression amount data in the same test, we make this claim with high confidence due to the almost same amount of impressions. and, at the end of the experiment phase, select the winner headline for all future impressions. This practice closely resembles the E&C algorithm in the bandit literature. We implement this method as a benchmark in our numerical experiments. The E&C method has a single hyperparameter to tune: the proportion of periods for exploration. We select the proportion from $\{0.1,0.2,0.3,0.4,0.5\}$ with the highest accumulated rewards. This ratio is fine-tuned for experiments with different total periods $T$, as a longer time horizon requires a smaller proportion of traffic for exploration, given a constant number of arms.
@@
Interestingly, the other two methods can be viewed as special cases of our proposed LLM-2UCBs algorithm. The Pure LLM (a.k.a pure exploitation, greedy algorithm) is a special case of LLM-2UCBs by setting $n^{\text{aux}}=\infty$, indicating complete trust in the LLM prediction results. This method fully relies on the LLM CTR prediction, directly selecting the headline with the highest predicted CTR throughout the entire horizon. Similarly, the standard UCB algorithm is a special case of LLM-2UCBs with $n^{\text{aux}}=0$, indicating no reliance on LLM CTR predictions. In our UCB implementation, we fine-tune the confidence control hyperparameter $\alpha$ from $\{0.02,0.04,\dots,0.10\}$ for different $T$. After fine-tuning, we find $\alpha=0.08$ in UCB gives the highest accumulated rewards. For LLM-2UCBs, we fine-tune the combination of hyperparameters $n^{\text{aux}}\in\{600,800,1000,1200,1400\}$ and $\alpha\in\{0.02,0.04,\dots,0.10\}$. After fine-tuning, we choose $n^{\text{aux}}=1000$ and $\alpha=0.08$ in LLM-2UCBs.
@@
After the fine-tuning of algorithms, we evaluate all algorithms under the test dataset, which has never been seen for CTR training and algorithm tuning. Algorithms do not know this true testing CTR in advance. We use the average CTR in the test data as the true CTRs to generate the click feedback in the numerical simulations. We argue that using this CTR as the true CTR in the simulation is justified because each headline was tested with a large amount of traffic, resulting in a relatively narrow confidence interval at Upworthy.
@@
UCB and LLM-2UCBs are both asymptotically optimal in minimizing regret, which means there is no difference between these two algorithms when the number of time horizons $T$ goes to infinity, and the LLM-based prior information is not helpful in the long run. However, in practice, the time horizon is never infinite. Therefore, we compare different algorithms under different time horizons. Since there can be a varying number of headlines/arms in different tests, we scale the time horizon using the metric “Traffic/Impression per headline” (we use the notion $\tau$ to represent) from $\{50,100,200,400,600,800,1000\}$. Specifically, if there are three headlines in one test and the multiplier is $100$, we set $T=100\times 3=300$ for this test. It is important to investigate the small number multipliers because they also partially address the question of what happens if we are concerned about inference and can only use a small amount of traffic for adaptive experiments.
@@ 4.3 Numerical Results
We visualize the numerical results in Figure 10 and document the pairwise comparison of different algorithms, reporting the relative percentage reward improvement and the p-value from the t-tests.
LLM-2UCBs generally performs the best among all algorithms. Notably, the Pure LLM approach exhibits consistent performance across different $\tau$ values because it is not an adaptive algorithm. For the other three approaches, which actively learn the CTR during the experiments, their average performance per period increases as the time horizon extends. This improvement is because the number of arms in the experiment remains constant, and with increasing periods, these algorithms have more opportunities to learn the CTRs, thus achieving better performance.
When comparing LLM-2UCBs with Pure LLM, we find that LLM-2UCBs outperforms Pure LLM in most cases. An exception occurs when $\tau=50$, where there is no significant difference between LLM-2UCBs and Pure LLM. This is due to the constrained exploration ability of LLM-2UCBs in short planning horizons, leading it to perform similarly to using only LLM prior information. However, as $\tau$ increases to 1000, LLM-2UCBs significantly outperforms Pure LLM, generating 5.46% more clicks. This observation indicates that we cannot fully rely on pure LLM predictions alone.
Comparing LLM-2UCBs with regular UCB, it is evident that leveraging LLM can always enhance UCB performance, although this improvement diminishes as the time horizon increases. This is due to the asymptotic optimality of the UCB algorithm, implying that as the time horizon approaches infinity, the initialization has a negligible effect on the average performance over an infinitely long period.
E&C consistently underperforms compared to LLM-2UCBs because it neither leverages LLM predictions nor achieves asymptotic optimality. Even when $\tau=1000$, LLM-2UCBs is still better than E&C by 4.41%, which is a large improvement in this industry. E&C is also significantly worse than Pure LLM when $\tau\leq 400$, and shows no significant difference from Pure LLM when $\tau\geq 600$. This suggests that the current Upworthy practice is suboptimal.
In summary, our findings underscore the superior performance of LLM-2UCBs, particularly as the time horizon increases, and highlight the limitations of relying solely on pure LLM predictions or the naive E&C approach.
@@
Parameter $\tau$ LLM-2UCBs vs Pure LLM LLM-2UCBs vs UCB LLM-2UCBs vs E&C Pure LLM vs UCB Pure LLM vs E&C UCB vs E&C
50 $0.53$ $4.06$ $6.63$∗∗∗∗ $3.52$ $6.07$∗∗∗ $2.46$
100 $1.04$ $5.01$∗∗∗∗ $6.20$∗∗∗∗ $3.93$∗∗ $5.11$∗∗∗∗ $1.13$
200 $1.92$∗∗∗∗ $4.14$∗∗∗∗ $4.89$∗∗∗∗ $2.18$ $2.92$∗∗ $0.72$
400 $3.57$∗∗∗∗ $2.51$∗∗∗ $6.36$∗∗∗∗ $-1.02$ $2.70$∗∗∗ $3.76$∗∗∗∗
600 $3.78$∗∗∗∗ $1.74$∗∗ $3.60$∗∗∗∗ $-1.96$∗∗ $-0.17$ $1.83$∗∗∗
800 $5.00$∗∗∗∗ $1.02$ $4.97$∗∗∗∗ $-3.79$∗∗∗∗ $-0.03$ $3.91$∗∗∗∗
1000 $5.46$∗∗∗∗ $1.10$ $4.41$∗∗∗∗ $-4.13$∗∗∗∗ $-1.00$ $3.27$∗∗∗∗
@@
Table 7: Percentage % reward improvement between two algorithms with significance levels: $p\leq$ 0.05, ∗∗ $p\leq$ 0.01, ∗∗∗ $p\leq$ 0.001, ∗∗∗∗ $p\leq$ 0.0001. For example, under the impression per headline equal to 50, the average click per test of LLM-2UCBs and Pure LLM is 0.05149 and 0.05122, respectively; so the relative improvement is calculated as ${0.05149}/{0.05122}-1=0.53\%$. Parameter $\tau$ represents the average impression per headline to scale the time horizon $T=\tau\times K$.

### v2

@@ 4.2 LOLA
In the first step, we use LLMs to get an initial prediction of each headline’s CTR. Note the prediction task here is somewhat different from that in §$\S$3. In the previous section, we used LLM-based approaches to perform pairwise comparisons of headlines and predict which headline amongst a pair is more engaging. However, as discussed earlier, that approach does not work when we have more than two headlines (which is the case for over 90% of the A/B tests in our data). Further, we need actual CTR predictions for each headline to feed into the second step of our algorithm. Therefore, we adapt the LLM-based classification models from §$\S$3 for predicting CTRs by making the following modifications: changing the loss function from log loss to mean squared error (MSE), using single headline inputs instead of headline pairs, and switching from binary output to continuous CTR prediction. The training and fine-tuning procedures remain unchanged. [^16] Using CTR as the outcome measure does not adversely affect the performance. For instance, using linear regression with 256-dimensional text embeddings from OpenAI to predict CTR for significant pairs and labeling the headline with the higher predicted CTR as the winner yields an accuracy of around 75%, which is very close to the 77.31% accuracy obtained with logistic regression. This minor difference is likely due to the inherently skewed nature of the CTR distribution and the sensitivity of MSE to outliers.
We compare two approaches for CTR prediction: an OpenAI-embedding-based method and a LoRA fine-tuned Llama-3-8b model. Our numerical results indicate that the R-squared values for the embedding-based approach and the LoRA fine-tuned method are 0.241 and 0.315, respectively. When selecting the headlines with the highest predicted CTRs, the accuracy—defined as the percentage of times the chosen headline indeed has the highest actual CTR—is 42.66% for the embedding-based method and 46.55% for the LoRA fine-tuned method. Note that this accuracy is significantly lower than the binary classification accuracy discussed in §$\S$3 because we can have more than two headlines in a test and many headlines have no significant difference in CTRs. Based on these results, we select the LoRA fine-tuned Llama-3-8b model as the best CTR prediction model in the first step. Additional details on the CTR prediction methods and their results can be found in Web Appendix §$\S$§F.
@@
In the second step, we build on the bandit framework from §$\S$4.1 to design an online algorithm that incorporates the LLM-based CTR predictions from the first step. Inspired by the 2-Upper Confidence Bounds (2-UCBs) policy designed by (gur2022adaptive), we propose Algorithm 1, titled LLM-Assisted 2-Upper Confidence Bounds (LLM-2UCBs). Essentially, in this algorithm, we treat the LLM’s CTR predictions as auxiliary information prior to the start of online experiments, equating this auxiliary CTR information to additional impressions and click outcomes for each arm.
@@
There are two hyperparameters in LLM-2UCBs – (1) The upper bound control hyperparameter, $\alpha$, that also shows up in the regular UCB algorithm, and (2) The hyperparameter $n^{\text{aux}}$, which is specific to our proposed algorithm and indicates the equivalent auxiliary samples for initialization. At a high level, we can regard the LLM predictions as prior information or the information flow before the algorithm starts. This prediction is approximately equivalent to initially generating $n^{\text{aux}}$ impressions for each arm and observing their outcomes to obtain the sample average CTR, $\bar{\mu}^{\text{aux}}_{k}$. The hyperparameter fine-tuning process is simple. We can utilize another randomly sampled dataset to test combinations of two hyperparameters and select the combination that yields the highest accumulated rewards.
@@
During the online learning phase, the key step is balancing the exploration and exploitation trade-off. We capture this trade-off using the 2-UCBs rule. Essentially, we have two valid UCBs: $U_{k}^{1}$ is the regular UCB, which uses observed impressions and clicks to construct the mean reward estimator and adds a regular upper bound. The second UCB, $U_{k}^{2}$, incorporates LLM predictions. The calculation of $U_{k}^{2}$ is straightforward. The mean reward component is simply the sample average, taking into account the “auxiliary” samples generated by the LLM before the learning phase begins. These auxiliary samples are imaginary and approximate the richness of information contained in the LLM prediction. The next step is to choose the smaller of the two UCBs, $\min\{U_{k}^{1},U_{k}^{2}\}$, as the final UCB for selecting the arm. Note that if $n^{\text{aux}}=0$, the 2-UCBs rule is the same as the standard UCB rule. Conversely, when $n^{\text{aux}}\rightarrow\infty$, the 2-UCBs rule is equivalent to a pure exploitation policy (a.k.a. greedy policy) using the LLM-based CTR prediction. Thus, when $n^{\text{aux}}$ is a positive integer, the 2-UCBs algorithm incorporates information from both LLMs and the actual click outcomes. After playing the arm and observing the outcome, we update the impressions, clicks, and mean reward accordingly, following standard bandit algorithm procedures.
The 2-UCBs algorithm offers three advantages. First, it is easy to implement, requiring the tuning of only two hyperparameters. Second, it can be used in conjunction with any LLM-based CTR prediction model in the first step and any outcome of interest in the second step. We use clicks as the outcome of interest given our application setting and data; however, other outcomes, such as time spent on a page or other measures of engagement, can be easily used instead of clicks. Finally, it is intuitive and easy to implement. Because it can easily accommodate the edge cases, e.g., one where the manager has no auxiliary information or one where the manager wants to fully rely on the LLM predictions, there is no need for extra engineering effort to set up separate systems for different edge cases. This adaptability is extremely valuable given the rapid advances in LLM technology. When the capability of the pure-LLM method continues to advance, the manager can easily fine-tune and increase $n^{\text{aux}}$ to take more advantage of LLM.
4.3 Benchmarks and Implementation Details
We now discuss the implementation details and present a set of benchmark algorithms. First, we randomly split our data into three exclusive folds based on the test ID, ensuring that all headlines in the same test are in the same fold and that the same headline does not appear in different folds. We use 70% of the data for training, 10% for hyperparameter fine-tuning, and 20% for testing. The training data is used to train the LLM models and is only used by our proposed LOLA approach (the LLM-2UCBs specifically) and pure LLM-based benchmarks. All hyperparameters in all the algorithms (both LOLA and the benchmark algorithms) are selected using the fine-tuning data, and the performance of the algorithms is evaluated on the test data.
We now describe a set of commonly used algorithms that serve as benchmarks and our LOLA approach.
@@ 4.3 Benchmarks and Implementation Details
    Explore-then-Commit (E&C): In this approach, the firm runs the A/B test for a fixed number of periods (i.e., explores) and then commits to the best-performing arm at that stage; see Chapter 6 (lattimore2020bandit). This is the standard A/B test approach used by many digital firms in the industry. In particular, Upworthy’s approach during the data collection period closely resembles the E&C algorithm – recall that they assign traffic to candidate headlines with equal probability during the experiment phase and then show the winning headline to all future impressions. The E&C method has a single hyper-parameter that needs tuning: the proportion of periods for exploration. We consider the following proportions – $\{0.1,0.2,0.3,0.4,0.5\}$, and choose the one with the highest accumulated rewards based on the 10% of data used for hyper-parameter tuning. We finally choose 0.2 for exploration.
@@
    Pure LLM-based approach: In this approach, the firm uses LLMs to predict the CTR for all the arms (headlines) and then allocates all the traffic to the arm with the highest predicted CTR. This is equivalent to a fully greedy policy based on LLM predictions. We use the training data to learn a model that predicts the CTR of a headline based on an LLM (either through text embeddings or LoRA fine-tuning). In our numerical study, we use the best-performing pure-LLM-based approach based on our experiments (see Web Appendix §$\S$§F), i.e., the LoRA fine-tuned Llama-3-8b with MSE loss, which gives a 46.55% accuracy. [^17] Note that since this is a pure prediction-based greedy approach, there are no hyper-parameters associated with the algorithm, and hence we do not use the fine-tuning data.
@@
    Standard UCB (UCB): This is the standard UCB algorithm discussed in §$\S$4.1. Here, the firm does not use any first-stage LLM predictions as an input, and there is only one hyperparameter to fine-tune, the confidence control, $\alpha$. We consider values from the set $\{0.02,0.04,\dots,0.10\}$, and find that $\alpha=0.08$ gives the highest accumulated rewards.
@@
    LOLA: This is our proposed approach. In the first step, we use CTR predictions from the best-performing pure-LLM-based CTR prediction model, i.e., the LoRA fine-tuning Llama-3-8b with MSE loss. Note that this is the same model that we use for the pure LLM-based approach, which allows for a fair comparison of the two approaches. For the second step, we use the LLM-2UCBs algorithm as outlined in §$\S$4.2. We fine-tune the combination of the two hyperparameters on the fine-tuning data, and we consider $n^{\text{aux}}\in\{600,800,1000,1200,1400\}$ and $\alpha\in\{0.02,0.04,\dots,0.10\}$. Based on fine-tuning, we choose $n^{\text{aux}}=1000$ and $\alpha=0.08$ for our setting.
@@
    Note that the former two methods can be viewed as special cases of our proposed LOLA algorithm. Both our algorithm and the pure LLM-based approach use the exact same CTR prediction algorithm. Hence, the pure LLM-based approach (a.k.a pure exploitation, greedy algorithm) is a special case of LOLA, wherein we set $n^{\text{aux}}=\infty$, indicating complete trust in the LLM prediction results. This method fully relies on the LLM CTR prediction, directly selecting the headline with the highest predicted CTR throughout the entire horizon. Similarly, the standard UCB algorithm is a special case of LOLA with $n^{\text{aux}}=0$, indicating no reliance on LLM CTR predictions. Again, note that we use the same hyper-parameter $\alpha=0.08$ across both algorithms, which ensures that our algorithm defaults to the standard UCB with $\alpha=0.08$ when we ignore the CTR predictions from the LLM.
@@
After the fine-tuning of algorithms, we evaluate all algorithms on the test dataset, which has never been used for either CTR training or algorithm tuning. Therefore, none of the algorithms know the true CTRs of the headlines in the test data in advance. We use the average CTR for each headline in the test data as the true CTRs to generate the click feedback in our numerical simulations. This approach of generating click outcomes is reasonable in our setting since each headline was tested on a large amount of traffic by Upworthy, resulting in relatively narrow confidence intervals.
@@
UCB and LOLA are both asymptotically optimal in minimizing regret, which means there is no difference between these two algorithms when the number of time horizons $T$ goes to infinity, and the LLM-based prior information is not helpful in an infinite horizon case. However, in practice, the time horizon is never infinite. Therefore, we compare all the algorithms for different values of time horizons. One challenge in our setting is that the number of headlines varies across tests. Intuitively, for any given time horizon, it is always easier to minimize regret (or identify the best headline) when the A/B test has only two headlines compared to a case when there are three headlines; and, of course, the problem only gets more challenging as the number of headlines increases. As a result, comparisons across time horizons are meaningful only when we keep the number of headlines constant. Since the number of headlines varies in our dataset (see Table 3), we scale the time horizon for each A/B test using the metric “Traffic/Impressions per headline” (which we represent using the notion $\tau$), where we choose $\tau$ from $\{50,100,200,400,600,800,1000\}$. For example, if there are three headlines in an A/B test and the multiplier is $\tau=100$, then we set $T=100\times 3=300$ for this test. Thus, the effective horizon across all tests for a given $\tau$ is the same. In our experiments, it will be particularly important to investigate the performance of different algorithms at small-medium multipliers because these represent situations where the media firm only has a small amount of traffic to work with and needs to adaptively experiment and effectively allocate impressions to the right headline with limited experimentation ability.
@@ 4.4 Numerical Results
We visualize the results from our numerical simulations in Figure 7 and document the pairwise comparison of different algorithms, along with the relative percentage reward improvement and the p-values from the t-tests in Table 8. We find that, on average, LOLA performs the best among all algorithms across all time horizons. We document a few additional patterns of interest.
Parameter $\tau$ LOLA vs UCB LOLA vs Pure LLM LOLA vs E&C UCB vs Pure LLM UCB vs E&C Pure LLM vs E&C
50 $6.95$∗∗∗∗ $0.09$ $8.69$∗∗∗∗ $-6.41$∗∗∗∗ $1.62$ $8.59$∗∗∗∗
100 $6.38$∗∗∗∗ $1.03$∗∗ $7.72$∗∗∗∗ $-5.02$∗∗∗∗ $1.26$ $6.61$∗∗∗∗
200 $4.98$∗∗∗∗ $1.78$∗∗∗∗ $8.16$∗∗∗∗ $-3.05$∗∗∗∗ $3.03$∗∗ $6.27$∗∗∗∗
400 $3.96$∗∗∗∗ $3.02$∗∗∗∗ $5.76$∗∗∗∗ $-0.90$ $1.74$ $2.66$∗∗∗
600 $3.04$∗∗∗∗ $3.49$∗∗∗∗ $4.83$∗∗∗∗ $0.43$ $1.73$∗∗ $1.30$
800 $2.35$∗∗∗∗ $4.20$∗∗∗∗ $5.36$∗∗∗∗ $1.81$∗∗∗ $2.94$∗∗∗∗ $1.12$
1000 $2.23$∗∗∗∗ $5.02$∗∗∗∗ $5.08$∗∗∗∗ $2.73$∗∗∗∗ $2.79$∗∗∗∗ $0.05$
@@
Table 8: Percentage % reward improvement between two algorithms with significance levels: $p\leq$ 0.05, ∗∗ $p\leq$ 0.01, ∗∗∗ $p\leq$ 0.001, ∗∗∗∗ $p\leq$ 0.0001. For example, under the impression per headline equal to 50, the average click per test of LOLA and UCB is 0.053760 and 0.050267, respectively; so the relative improvement is calculated as ${0.053760}/{0.050267}-1=6.95\%$. Parameter $\tau$ represents the average impression per headline to scale the time horizon $T=\tau\times K$.
First, the Pure LLM approach exhibits consistent performance across different $\tau$ values because it is not an adaptive algorithm. The small fluctuation at $\tau\in\{50,100,200\}$ is due to the randomness of average clicks per period, which tends to be larger under the smaller time horizon. For the other three approaches, which actively learn the CTR during the experiments, their average performance per period increases as the time horizon extends. This improvement is because the number of arms in the experiment remains constant, and with increasing periods, these algorithms have more opportunities to learn the CTRs, thus achieving better performance. Comparing the performance of LOLA with pure LLM methods, we find that LOLA generally outperforms Pure LLM, except when $\tau=50$, where there is no significant difference between LOLA and Pure LLM. This is because the short planning horizon does not allow LOLA to learn new information over and above the information already available from the LLM. As $\tau$ increases to 1000, we see that LOLA significantly outperforms Pure LLM, generating 5.02% more clicks. This suggests that if the horizon is very short, the planner/firm will not see significant improvement over simply relying on LLM predictions. However, as the horizon increases, LOLA will start showing significant gains.
Next, comparing LOLA with regular UCB, we see that it always outperforms UCB. Interestingly, here we see that, as the time horizon increases, the relative improvement from LOLA diminishes. Intuitively, this is because both UCB and LOLA are active learning algorithms, and as the time horizon increases, they are both able to learn the CTRs better. While LOLA starts with a relative information advantage, as the time horizon increases, this advantage diminishes. Formally, this pattern is due to the asymptotic optimality of the UCB and LOLA algorithms, implying that as the time horizon approaches infinity, the initialization has a negligible effect on the average performance over an infinitely long period. We also see that E&C consistently underperforms compared to LOLA because it neither leverages LLM predictions nor achieves asymptotic optimality. While E&C shows better performance as the horizon increases, even when $\tau=1000$, LOLA is still better than E&C by 5.08%, which is a large improvement in the media and publishing industry.
Finally, comparing the benchmark algorithms with each other, we see that when the horizon is short, the Pure LLM does better, while UCB is better when the horizon is longer. In general, this reflects the intuition that active learning is not very useful when the time horizons are very short since the algorithm has no time to explore and learn the CTRs effectively (and then exploit this learning). So simply going with the pure LLM predictions is better. In contrast, when the horizons are longer, UCB (or active learning) outperforms using static predictions from the LLM. Note that LOLA combines the strengths of both UCB and Pure LLM, and hence it is able to outperform both and provide superior performance in both short and long horizons. We also note that E&C is generally worse than both Pure LLM and UCB, which suggests that the current Upworthy practice is suboptimal. This is because E&C neither exploits LLM information and nor does it actively learn over time.
In summary, we find that LOLA outperforms existing approaches for experimentation by leveraging the strengths of LLMs and combining them with the benefits of active/online learning. In particular, our results demonstrate the value of LOLA as the time horizon increases, and highlight the limitations of relying solely on pure LLM predictions or the naive E&C approach.
4.5 LOLA Variants and Extensions
So far, we have considered a version of LOLA that uses a fine-tuned CTR prediction model for LLM predictions in the first stage and an LLM-2UCBs algorithm that minimizes regret for online learning in the second step. However, LOLA is a general framework and can be easily adapted to a wide variety of settings. We highlight a few natural extensions and variants below. Details of the models and numerical results (when applicable) are shown in Web Appendix §$\S$§G.
Best Arm Identification: So far, we focused on the goal of regret minimization, where the goal is to maximize the clicks/reward (per Equation (1)). While regret minimization is the most commonly studied goal in active learning and the natural goal for businesses in most settings, another commonly studied goal is Best Arm Identification (BAI). The goal of BAI problems is to identify the arm with the highest reward as fast and accurately as possible, regardless of accumulated regret. Formally, given a low failure rate $\delta$, we aim to find the arm $a^{*}$ with the highest expected reward, at least $1-\delta$ probability, while minimizing the number of total pulls. [^18] This is also called fixed-confidence setting (garivier2016optimal). Another common approach is the fixed-budget setting, where the goal is to maximize the probability of finding the best arm within a fixed number of pulls. We can easily modify the LOLA framework for a BAI goal. In Web Appendix §$\S$§G.1, we present a LLM-BAI algorithm that builds on the recently proposed BAI algorithm by mason2020finding. This algorithm is particularly effective in our setting since it can accommodate scenarios where the difference between the rewards of arms is small or insignificant. We show that our LLM-BAI algorithm significantly outperforms both the plain BAI algorithm (proposed by mason2020finding) as well as the standard A/B test under the same failure rate restriction. Please see Web Appendix §$\S$§G.1 for a detailed description of the LLM-BAI algorithm and its performance against the standard benchmarks.
Thompson Sampling: Our bandit specification in §$\S$4.1 reflects a stochastic bandit setting rather than a Bayesian bandit setting. However, our proposed algorithms can easily be applied in a Bayesian setting. We present a Thompson sampling version of our proposed LOLA algorithm (LLM-TS) and its numerical performance in Web Appendix §$\S$§G.2. We see that UCB-based algorithms perform better than Thompson Sampling-based algorithms in both the LLM-assisted version and the standard version. That is, LLM-TS performs worse than LLM-2UCBs, and standard TS performs worse than UCB. This pattern has been observed in the literature, and it is generally known that TS suffers from over-exploration (min2020policy); in other words, UCB’s upper confidence shrinks faster, leading it to exploit more effectively. Nevertheless, in business settings where the firm already has a TS-based adaptive experimentation framework, it is easy to adapt the LOLA approach within this approach.
User Information: So far, we have not considered user-level features or how to personalize the content shown to different users based on their behavioral/context features (mostly because the UpWorthy data does not have user-level features). Prior research has shown that using such features can significantly improve the match between users and content (li2010contextual; yoganarasimhan_2020; rafieian_yoganarasimhan_2021). It is, however, easy to expand our LOLA approach to include user features in both the LLM training phase as well as the online learning phase. We present two versions of LOLA that extend existing contextual bandit algorithms to our setting. First, we can extend the standard contextual linear bandits (chu2011contextual) to our LOLA framework by incorporating both user and text features in the LLM training and online learning phases of LOLA. However, this linear bandit approach has limitations since it assumes that the rewards model is linear in text and user features. Therefore, we also propose another solution based on the FALCON algorithm that does not make any linearity assumption (simchi2022bypassing). We present the details of both these contextual LOLA algorithms and their pros and cons in Web Appendix §$\S$§G.3.
Alternative rewards: In the Upworthy setting, the measure of reward is clicks. However, click-based metrics have raised concerns about promoting clickbait headlines, which may reduce user engagement or spread negativity. As a result, the news and media industry is now evolving to focus more on content quality and humanity, although clicks still remain an important measure; see (Upworthy2023) for a more detailed discussion. The LOLA framework can be easily adapted to alternative reward measures that align with new business goals, e.g., time spent on the platform during a session, or time spent on the article (and not just clicks on the headline).
Apart from these extensions, the LOLA approach is quite general and adaptable on other dimensions as well. For instance, it is agnostic to the exact nature of the LLM-based approach used in the first step. Any of the LLM-based approaches discussed in §$\S$3 can be used in the first step. This is important since the LLM models and fine-tuning approaches continue to advance quickly. Further, LOLA can also be used in conjunction with content/creatives created by LLMs themselves. For instance, one simple way to leverage fine-tuned models is to use them to generate content (instead of using human-generated headlines/content). Indeed, recent studies have shown that promotional content (ads/emails) generated by fine-tuned LLMs tend to outperform traditional personalized ads and human-generated content in effectiveness (kumar2023generative; angelopoulos2024value). LLM-generated headlines/content can be used as competing arms/treatments within our LOLA framework. Additionally, the LOLA framework can also be easily extended to content recommendation systems used by social media platforms such as X and Facebook. In summary, the plug-and-play nature of our approach can easily accommodate newer advances and developments at all stages of the implementation process.